OpenStudy (anonymous):

Write the equation for 5x + 2y = 3 in slope-intercept form.

OpenStudy (lifeisadangerousgame):

Okay so do you know what slope-intercept form is?

OpenStudy (anonymous):

y=5/2x+3/2

OpenStudy (anonymous):

You basically just want to isolate y in slope intercept form.

OpenStudy (lifeisadangerousgame):

Slope intercept form is y = mx + b So for the equation 5x + 2y = 3 We need to get the y on the left by itself and the 5x on the other side so its these steps: 5x + 2y = 3 -5x -5x 2y = -5x + 3 divide by 2 on both sides y = 5/2x + 3/2 That should be the answer, let me check

OpenStudy (anonymous):

You're good life

OpenStudy (lifeisadangerousgame):

Haha Slope-intercept is my favorite form so I know it well :3

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I try to be the Jack of all trades :3

OpenStudy (anonymous):

I Hate Slope Interform Dats What I HAve Troube WIth All DA Time! Thank You!

OpenStudy (anonymous):

n p

OpenStudy (lifeisadangerousgame):

Slope intercept is the easiest form, believe you me (for me anyway) xD